# Planetary Edge Trends (PET)

### v1

@@ Abstract
Context. Recent advancements in exoplanet detection have led to over 5,700 confirmed detections. The planetary systems hosting these exoplanets exhibit remarkable diversity.
@@
Methods. In this work, based on the $Kepler$ Data Release 25 (DR25) catalog combined with LAMOST and $Gaia$ data, we investigate the correlation between stellar mass and the inner edge position across different populations of small planets in multi-planetary systems, such as super-Earths and sub-Neptunes. By correcting for the influence of stellar metallicity and analyzing the impact of observational selection effects, we confirm the trend that as stellar mass increases, the position of the inner edge shifts outward.
Results. Our results reveal a stronger correlation between the inner edge and stellar mass ($a_{\text{in}}\propto M_{\star}^{\gamma_{1}}$) with a power-law index of $\gamma_{1}=0.6-1.1$, which is larger compared to previous studies. The stronger correlation in our findings is primarily attributed to two factors: first, the metallicity correction applied in this work enhances the correlation; second, the previous use of occurrence rates to trace the inner edge weakens the observed correlation.
Conclusions. Through comparison between observed statistical results and current theoretical models, we find that the pre-main-sequence (PMS) dust sublimation radius of the protoplanetary disk best matches the observed inner edgestellar mass. Therefore, we conclude that the inner dust disk likely limits the innermost orbits of small planets, contrasting with the inner edges of hot Jupiters, which are associated with the magnetospheres of gas disks, as suggested by previous studies. This highlights that the inner edges of different planetary populations are likely regulated by distinct mechanisms.
@@ 1 Introduction
The emergence of exoplanet surveys, such as the $Kepler$ mission, has revolutionized our understanding of the diversity and abundance of planetary systems (Borucki, 2016). With the advancement of exoplanet detection, over 5,700 exoplanets have been confirmed to date (https://exoplanetarchive.ipac.caltech.edu/index.html), and these exoplanets reside in planetary systems that exhibit remarkable diversity (Zhu & Dong, 2021). The uniqueness of our Solar System has long been a topic of interest (Zhu et al., 2018). For instance, Mercury, the innermost planet in our Solar System, orbits the Sun at a distance of about 0.4 AU. In contrast, many planetary systems host planets that orbit much closer to their host stars, within 0.4 AU (Mulders et al., 2018). This raises the question: is our Solar System a peculiar case among these diverse planetary systems? To address this, we aim to investigate the intrinsic patterns reflected by the inner edges of planetary systems, specifically focusing on the positions of the innermost planets. Gaining a deep understanding of the intricacies of where these planets reside within their respective systems (Millholland et al., 2022; Sobski & Millholland, 2023) provides invaluable insights into the fundamental processes that govern planetary formation, migration, and stability. In recent years, a wealth of observational and theoretical research has illuminated the complex relations between stellar properties and planetary architectures (Winn & Fabrycky, 2015; Zhu & Dong, 2021; Weiss et al., 2023). Therefore, studying the relationship between the distribution of the inner edge and stellar properties is crucial.
@@
The inner edge of protoplanetary disks around pre-main-sequence solar-type stars is located at approximately 0.1 AU (Millan-Gabet et al., 2007), which corresponds with findings in several statistical studies that observed clustering of the inner edges of planetary systems around a 10-day (0.1 AU) orbital period (Lee & Chiang, 2017; Mulders et al., 2018). Some theoretical studies have analyzed the mechanisms that determine the inner edge. For in-situ formation, the location of the inner edge of the protoplanetary disk is related to the dust sublimation radius (Dullemond et al., 2001; Pinte et al., 2008; Kama et al., 2009; Min et al., 2011; Tripathi et al., 2017; Liu et al., 2019). Due to the lack of solid material required for in-situ formation (Boley & Ford, 2013), planets are less likely to be formed within these radii (Chiang & Laughlin, 2013). For disk migration, planets form at more distant locations and migrate inward through the gas disk, becoming trapped at the co-rotation radius or the sublimation radius (Lin et al., 1996; Masset et al., 2006; Kley & Nelson, 2012; Swift et al., 2013; Batygin et al., 2023). As a result, it is less likely for planets to exist within these radii (Kuchner & Lecar, 2002). Stellar tides and planetary tides within multi-planetary systems can also affect the radii where there is a lack of planets (Wu et al., 2007; Jackson et al., 2009). However, it remains unclear which theoretical model plays a dominant role, and further observational constraints and testing are needed to determine the most plausible model. In Table 2, we summarize the theoretical models discussed above.
@@
As observational data continue to increase, using large datasets like those from $Kepler$ to test inner-edge theoretical models has become possible. Several observational studies have investigated the relationship between the inner edge of planetary systems and stellar mass using a power-law model: $a_{\text{in}}\propto M_{\star}^{\gamma_{1}}$, where the power-law index $\gamma_{1}$ represents this correlation (as in our work). Plavchan & Bilinski (2013) analyzed exoplanets orbiting within 0.1 AU, including confirmed Jovian exoplanets and $Kepler$ planetary candidates from the early data release. Their findings confirmed a correlation between the planetary migration halting distance (i.e., the inner edge) and stellar mass, which they examined through this power-law model. They found that migration halting from tidal circularization mechanism (Ford & Rasio, 2006; Wu et al., 2007) provided the best fit for the empirical distribution of confirmed Jovian exoplanets. However, for $Kepler$ candidates, they derived a power-law index of $\gamma_{1}=0.38-0.9$, which exceeds the tidal halting mechanism’s prediction of $\gamma_{1}=0.23-0.33$. Mulders et al. (2015) analyzed the entire $Kepler$ sample of smaller planets to assess planet occurrence around stars of varying spectral types, revealing notable differences across these types. The occurrence drop depended on stellar mass and scaled with the semi-major axis with a power-law index of $\gamma_{1}\sim 0.33$. Their findings indicated that this scaling closely aligned with the pre-main-sequence co-rotation radius, leading them to derive a smaller power-law index for the planet occurrencestellar mass scaling law. Mendigutía et al. (2024), by combining TESS and $Gaia$ DR3 data, identified a sample of 47 intermediate-mass main-sequence stars hosting confirmed and strong candidate hot Jupiters. Their findings indicated that hot Jupiters around intermediate-mass stars tended to orbit closer to the central stars than the inner dust disk, generally aligning with the magnetospheric truncation radius. Their result suggested that the inner gas disk, rather than the dust disk, constrained the innermost orbits of hot Jupiters around intermediate-mass stars.
@@
In this paper, to address these issues, we systematically analyze the relationship between the inner edge of planetary systems and stellar mass leveraging comprehensive and enriched observational data from LAMOST (Chen et al., 2021a, b), $Gaia$ (Berger et al., 2020), and $Kepler$ (Thompson et al., 2018). By combining LAMOST’s spectroscopic measurements with $Gaia$’s astrometric data, we achieve precise characterization of $Kepler$ host stars, offering a large and homogeneous sample with accurate stellar properties. For example, the uncertainty in stellar mass is only 7% (Berger et al., 2020). Furthermore, LAMOST provides a large sample of host stars with metallicity measurements (Dong et al., 2014), which we utilize to investigate the impact of stellar metallicity on the inner edgestellar mass relationship. Given that metallicity may influence the inner edge (Mulders et al., 2016) and is highly correlated with stellar mass (Johnson et al., 2010), understanding its impact on the dependence of the inner edge on stellar mass is essential. Additionally, motivated by the radius gap observed in $Kepler$ systems (Owen & Wu, 2013; Fulton et al., 2017), we further divide planetary systems into super-Earth, sub-Neptune, and mixed systems for a more detailed analysis to better reflect the relationship between the inner edge and stellar mass.
This work is organized as follows: In Sect. 2, we introduce the samples used in this study and the process of selecting both stellar and planetary samples. Then, in Sect. 3, we present the results of our statistical analyses to explore the observational findings. In Sect. 4, we compare and discuss our findings with theoretical models, extend our analysis by comparing them with previous studies, and provide insights for future work. Finally, in Sect. 5, we summarize the results and conclusions of this study and provide perspectives for future research. Additional detailed discussions and analyses are included in the Appendix A.
@@ 2.1 Dataset
The planetary properties are sourced from the $Kepler$ Data Release 25 (DR25) catalog (Thompson et al., 2018), which includes 8054 $Kepler$ Objects of Interest (KOIs). The primary stellar properties we analyzed, such as stellar mass, are based on the Berger et al. (2020) $Gaia$-$Kepler$ Stellar Properties Catalog (hereafter referred to as Berger20), which includes 186,301 stars. Stellar metallicity data is derived from PAST II (Chen et al., 2021b), utilizing the LAMOST-$Gaia$-$Kepler$ Catalog (hereafter referred to as PAST II), which contains 35,835 stars and 1,060 planets. Systems lacking metallicity data (1,294 stars) are removed from our analysis. By combining data from $Kepler$ DR25, Berger20, and PAST II, we establish the dataset for this study as shown in Table 1, which serves as the standard dataset in this paper.
Table 1: Data Selection
@@ 2.2 Stars and Planets
We use the data filtering process from Table 1 as the standard case and select stars and planets according to the following filters:
@@
    Not a false positive. We exclude planets flagged as false positives from our analysis.
@@
    Not a binary star. We use RUWE $<$ 1.2 to exclude binary stars (Berger et al., 2020). RUWE, the Renormalized Unit Weighted Error, provides a renormalized metric that is independent of color and magnitude, helping to identify and exclude binary systems.
@@
    Not a giant star. We exclude giant stars using the criterion $\log_{10}{\frac{\mathrm{R_{\star}}}{\mathrm{R_{\odot}}}}<0.00035\times(\mathrm{T_{\rm eff}}-4500)+0.15$ as outlined in Fulton et al. (2017), focusing our study on main-sequence systems.
@@
    Planet orbital period $(\rm P)$. We exclude planet candidates with orbital periods shorter than 1 day, as ultra-short-period (USP) planets represent a distinct planetary population (Dai et al., 2018).
@@
    Planet radius $(\rm R_{\rm p})$. We exclude giant planets (with radii greater than 4 $\rm R_{\oplus}$) and primarily focus on small-radius planets.
@@
    Detection efficiency. We consider planets with a signal-to-noise ratio (SNR) greater than 7.1, as this is the detection threshold calculated by the $Kepler$ team (Borucki et al., 2011).
After applying the above selection criteria, we are left with 587 stars hosting a total of 829 planets. We categorize these planetary systems into multiple-transiting planet systems (hereafter multis) and single-transiting planet systems (hereafter singles), with 166 and 421 host stars, respectively. However, this study focuses exclusively on multis. Fig. 1 shows the distribution of stars and planets analyzed in this paper.
In $Kepler$ systems, singles constitute the vast majority. Due to the fact that singles tend to have larger inclinations (Xie et al., 2016; Zhu et al., 2018), the innermost planet (the true inner edge) may be missed, resulting in the observation of only the outer planets (a false inner edge), which introduces observational bias in measuring the inner edge. In contrast, multis in $Kepler$ systems are generally considered to be coplanar (Fabrycky et al., 2014; Xie et al., 2016), making it unlikely to detect outer planets without also observing the innermost ones. However, this does not apply to ultra-short-period (USP) planets (Dai et al., 2018), which have already been excluded from our sample. Therefore, we separate singles and multis, conducting independent studies on the latter. We believe that the observational results from multis are more representative of the intrinsic relationship between the inner edge and stellar properties.
@@ 2.3 Systems Division
Figure 2: The architecture of the three population samples of all multis studied in this paper: the left panel represents super-Earths, the middle panel represents sub-Neptunes, and the right panel represents mixed systems containing both super-Earths and sub-Neptunes within the same system. The point colors and styles are the same as those in Fig. 1. Each row represents a system, ordered from top to bottom by decreasing stellar mass. A trend can be observed: as stellar mass increases from low to high, the semi-major axis shifts from smaller to larger values.
Previous studies have revealed a valley around 1.9 R in the radius distribution of small planets, which separates super-Earths and sub-Neptunes (Owen & Wu, 2013; Fulton et al., 2017). To facilitate a more detailed analysis, we calculate the radius valley using Eq. (1):
@@
which is derived from Eq. (12) in Zhu & Dong (2021). Based on this, we divide all planetary systems in our sample into three categories: systems containing only super-Earths (Rp $<$ Rgap), systems containing only sub-Neptunes (Rp $\geq$ Rgap), and mixed systems that include both super-Earths and sub-Neptunes within the same system. These three categories represent completely distinct populations with no overlap.
Fig. 2 illustrates the architecture of the three population samples of all multis analyzed in this study. The left panel represents super-Earth systems, the middle panel represents sub-Neptune systems, and the right panel represents mixed systems containing both super-Earths and sub-Neptunes within the same system. Each panel is sorted by stellar mass. A trend is evident: as stellar mass increases from low to high, the semi-major axis shifts progressively from smaller to larger values.
In the final sample, we use the following data: All Multiple Systems consist of 166 systems with 408 planets; Super-Earth Systems include 55 systems with 132 planets; Sub-Neptune Systems contain 33 systems with 70 planets; and Mixed Multiple Systems comprise 78 systems with 206 planets. The “All Multiple Systems” dataset includes the total data from the other three populations.

### v2

@@ Abstract
Context. Recent advancements in exoplanet detection have led to over 5,900 confirmed detections. The planetary systems hosting these exoplanets exhibit remarkable diversity.
@@
Methods. In this work, based on the $Kepler$ Data Release 25 catalog combined with LAMOST and $Gaia$ data, we investigate the correlation between stellar mass and the inner edge position across different populations of small planets in multi-planetary systems, such as super-Earths and sub-Neptunes. By correcting for the influence of stellar metallicity and analyzing the impact of observational selection effects, we confirm the trend that as stellar mass increases, the position of the inner edge shifts outward.
Results. Our results reveal a stronger correlation between the inner edge and stellar mass ($a_{\text{in}}\propto M_{\star}^{\gamma_{1}}$), with a power-law index of $\gamma_{1}=0.6-1.1$, which is larger compared to previous studies. The stronger correlation in our findings is primarily attributed to two factors: first, the metallicity correction applied in this work enhances the correlation; second, the previous use of occurrence rates to trace the inner edge weakens the observed correlation.
Conclusions. Through comparison between observed statistical results and current theoretical models, we find that the pre-main-sequence dust sublimation radius of the protoplanetary disk best matches the observed inner edgestellar mass. Therefore, we conclude that the inner dust disk likely limits the innermost orbits of small planets, contrasting with the inner edges of hot Jupiters, which are associated with the magnetospheres of gas disks, as suggested by previous studies. This highlights that the inner edges of different planetary populations are likely regulated by distinct mechanisms.
@@ 1 Introduction
The emergence of exoplanet surveys, such as the $Kepler$ mission, has revolutionized our understanding of the diversity and abundance of planetary systems (Borucki, 2016). With the advancement of exoplanet detection, over 5,900 exoplanets have been confirmed to date, [^1] https://exoplanetarchive.ipac.caltech.edu/index.html and these exoplanets reside in planetary systems that exhibit a remarkable diversity (Zhu & Dong, 2021). The uniqueness of our Solar System has long been a topic of interest (Zhu et al., 2018). For instance, Mercury, the innermost planet in the Solar System, orbits the Sun at a distance of about 0.4 AU. In contrast, many planetary systems host planets that orbit much closer to their host stars, within 0.4 AU (Mulders et al., 2018). Given the diverse planetary systems observed, an important question arises regarding how our own Solar System should be explored, understood, and interpreted. To address this, we aim to investigate the intrinsic patterns reflected by the inner edges of planetary systems. Specifically, we focus on the positions of the innermost planets. Gaining a deep understanding of the intricacies of where these planets reside within their respective systems (Millholland et al., 2022; Sobski & Millholland, 2023) will provide invaluable insights into the fundamental processes that govern planetary formation, migration, and stability. In recent years, a wealth of observational and theoretical research has illuminated the complex relations between stellar properties and planetary architectures (Winn & Fabrycky, 2015; Zhu & Dong, 2021; Weiss et al., 2023). Therefore, studying the relationship between the distribution of the inner edge and stellar properties is crucial.
@@
The inner edge of protoplanetary disks around pre-main-sequence solar-type stars is located at approximately 0.1 AU (Millan-Gabet et al., 2007), which corresponds with findings in several statistical studies that observed clustering of the inner edges of planetary systems around a 10-day (0.1 AU) orbital period (Lee & Chiang, 2017; Mulders et al., 2018). Some theoretical studies have analyzed the mechanisms that determine the inner edge. For in situ formation, the location of the inner edge of the protoplanetary disk is related to the dust sublimation radius (Dullemond et al., 2001; Pinte et al., 2008; Kama et al., 2009; Min et al., 2011; Tripathi et al., 2017; Liu et al., 2019). Due to the lack of solid material required for in situ formation (Boley & Ford, 2013), planets are less likely to be formed within these radii (Chiang & Laughlin, 2013). For disk migration, planets form at more distant locations and migrate inward through the gas disk, becoming trapped at the co-rotation radius or the sublimation radius (Lin et al., 1996; Masset et al., 2006; Kley & Nelson, 2012; Swift et al., 2013; Batygin et al., 2023). As a result, it is less likely for planets to exist within these radii (Kuchner & Lecar, 2002). Stellar tides and planetary tides within multi-planetary systems can also affect the radii where there is a lack of planets (Wu et al., 2007; Jackson et al., 2009). However, it remains unclear which theoretical model plays a dominant role, and further observational constraints and testing are needed to determine the most plausible model. In this work, we summarize the theoretical models discussed above.
@@
As observational data continue to increase, using large datasets such as those from $Kepler$ to test inner-edge theoretical models has become possible. Several observational studies have investigated the relationship between the inner edge of planetary systems and stellar mass using a power-law model: $a_{\text{in}}\propto M_{\star}^{\gamma_{1}}$, where the power-law index $\gamma_{1}$ represents this correlation (as in our work). Plavchan & Bilinski (2013) analyzed exoplanets orbiting within 0.1 AU, including confirmed Jovian exoplanets and $Kepler$ planetary candidates from the early data release (i.e., the third tabulation of $Kepler$ planetary candidates). Their findings confirmed a correlation between the planetary migration halting distance (i.e., the inner edge) and stellar mass, which they examined through this power-law model. They found that migration halting from tidal circularization mechanism (Ford & Rasio, 2006; Wu et al., 2007) provided the best fit for the empirical distribution of confirmed Jovian exoplanets. However, for $Kepler$ candidates, they derived a power-law index of $\gamma_{1}=0.38-0.9$, which exceeds the tidal halting mechanism’s prediction of $\gamma_{1}=0.23-0.33$. Mulders et al. (2015) analyzed the entire $Kepler$ sample of smaller planets to assess planet occurrence around stars of varying spectral types, revealing notable differences across these types. The occurrence drop depended on stellar mass and scaled with the semimajor axis with a power-law index of $\gamma_{1}\sim 0.33$. Their findings indicated that this scaling closely aligned with the pre-main-sequence co-rotation radius, leading them to derive a smaller power-law index for the planet occurrencestellar mass scaling law. Mendigutía et al. (2024), by combining TESS and $Gaia$ DR3 data, identified a sample of 47 intermediate-mass main-sequence stars hosting confirmed and strong candidate hot Jupiters. Their findings indicated that hot Jupiters around intermediate-mass stars tended to orbit closer to the central stars than the inner dust disk, generally aligning with the magnetospheric truncation radius. Their result suggested that the inner gas disk, rather than the dust disk, constrained the innermost orbits of hot Jupiters around intermediate-mass stars.
@@
In this paper, to address these issues, we systematically analyze the relationship between the inner edge of planetary systems and the stellar mass by leveraging comprehensive and enriched observational data from LAMOST (Chen et al., 2021a, b), $Gaia$ (Berger et al., 2020), and $Kepler$ (Thompson et al., 2018). By combining LAMOST’s spectroscopic measurements with $Gaia$’s astrometric data, we are able to achieve precise characterization of $Kepler$ host stars, offering a large and homogeneous sample with accurate stellar properties. For example, the uncertainty in stellar mass is only 7% (Berger et al., 2020). Although this is the best available stellar mass data for this study, it is important to note that the 7% refers to precision rather than accuracy. Furthermore, LAMOST provides a large sample of host stars with metallicity measurements (Dong et al., 2014), which we utilize to investigate the impact of stellar metallicity on the inner edgestellar mass relationship. Given that metallicity may influence the inner edge (Mulders et al., 2016) and is highly correlated with stellar mass (Johnson et al., 2010), understanding its impact on the dependence of the inner edge on stellar mass is essential. Additionally, motivated by the radius gap observed in $Kepler$ systems (Owen & Wu, 2013; Fulton et al., 2017), we further divide planetary systems into super-Earth, sub-Neptune, and mixed systems for a more detailed analysis and to better reflect the relationship between the inner edge and stellar mass.
This work is organized as follows: In Sect. 2, we introduce the samples used in this study and the process of selecting both stellar and planetary samples. Then, in Sect. 3, we present the results of our statistical analyses to explore the observational findings. In Sect. 4, we compare and discuss our findings with theoretical models, extend our analysis by comparing our results with previous studies, and provide insights for future work. Finally, in Sect. 5, we summarize the results and conclusions of this study and provide perspectives for future research. Additional detailed discussions and analyses are included in the appendix A.
@@ 2.1 Dataset
The planetary properties are sourced from the $Kepler$ Data Release 25 (DR25) catalog (Thompson et al., 2018), which includes 8,054 $Kepler$ Objects of Interest (KOIs). The primary stellar properties we analyzed, such as stellar mass, are based on the Berger et al. (2020) $Gaia$-$Kepler$ Stellar Properties Catalog (hereafter referred to as Berger20), which includes 186,301 stars. Stellar metallicity data is derived from PAST II (Chen et al., 2021b), utilizing the LAMOST-$Gaia$-$Kepler$ Catalog (hereafter referred to as PAST II), which contains 35,835 stars and 1,060 planets. Systems lacking metallicity data (1,294 stars) were removed from our analysis. By combining data from $Kepler$ DR25, Berger20, and PAST II, we established the dataset for this study as shown in Table 1.
Table 1: Data selection.
@@ 2.2 Stars and planets
We used the data filtering process from Table 1 as the standard case and selected stars and planets according to the following filters:
@@
    Not a false positive. We excluded planets flagged as false positives from our analysis.
@@
    Not a binary star. The renormalized unit weighted error (RUWE) provides a renormalized metric that is independent of color and magnitude, and it helps identify and exclude binary systems. Therefore, we used RUWE $<$ 1.2 to exclude binary stars (Berger et al., 2020).
@@
    Not a giant star. We excluded giant stars using the criterion $\log_{10}{\frac{\mathrm{R_{\star}}}{\mathrm{R_{\odot}}}}<0.00035\times(\mathrm{T_{\rm eff}}-4500)+0.15$ as outlined in Fulton et al. (2017) in order to focus our study on main-sequence systems.
@@
    Planet orbital period $(\rm P)$. We excluded planet candidates with orbital periods shorter than one day, as ultra-short-period planets represent a distinct planetary population (Dai et al., 2018).
@@
    Planet radius $(\rm R_{\rm p})$. We excluded giant planets (i.e., planets with radii greater than 4 $\rm R_{\oplus}$) and primarily focused on small-radius planets.
@@
    Detection efficiency. We considered planets with a signal-to-noise ratio (S/N) greater than 7.1, as this was the detection threshold calculated by the $Kepler$ team (Borucki et al., 2011).
After applying the above selection criteria, we were left with 587 stars hosting a total of 829 planets. We categorized these planetary systems into multiple-transiting planet systems (hereafter multis) and single-transiting planet systems (hereafter singles), with 166 and 421 host stars, respectively. However, this study focused exclusively on multis. Figure 1 shows the distribution of stars and planets analyzed in this paper.
During the selection criteria process, we considered stellar multiplicity and the host star’s evolutionary stage as potential factors that could influence the inner edge. To obtain a cleaner sample, we aimed to isolate the effect of stellar mass as much as possible while minimizing the impact of other factors. The formation and evolution of planetary systems around binary stars may differ significantly from those around single stars (e.g., Duchêne & Kraus, 2013). Binaries can alter the structure of the protoplanetary disk and affect planetary orbits (Kraus et al., 2016; Christian et al., 2022; Dupuy et al., 2022), thereby influencing the position of the inner edge. Moreover, the impact of the host star’s evolutionary stage on planetary orbits is more direct. As a star evolves, its temperature changes, and when it transitions into a red giant, these changes may lead to the engulfment of the innermost planets.
Figure 2: Architecture of the three population samples of all multis studied in this paper. The left panel represents super-Earths, the middle panel represents sub-Neptunes, and the right panel represents mixed systems containing both super-Earths and sub-Neptunes within the same system. The point colors and styles are the same as those in Fig. 1. Each row represents a system, ordered from top to bottom by decreasing stellar mass. A trend can be observed: As stellar mass increases from low to high, the semimajor axis shifts from smaller to larger values.
In $Kepler$ systems, singles constitute the vast majority. Due to the fact that singles tend to have larger inclinations (Xie et al., 2016; Zhu et al., 2018), the innermost planet (the true inner edge) may be missed, resulting in the observation of only the outer planets (a false inner edge), which introduces observational bias in measuring the inner edge. In contrast, multis in $Kepler$ systems are generally considered to be coplanar (Fabrycky et al., 2014; Xie et al., 2016), making it unlikely to detect outer planets without also observing the innermost ones. However, this does not apply to ultra-short-period planets (Dai et al., 2018), which have already been excluded from our sample. Therefore, we separated singles and multis, conducting independent studies on the latter. We believe that the observational results from multis are more representative of the intrinsic relationship between the inner edge and stellar properties.
@@ 2.3 Systems division
Previous studies have revealed a valley around 1.9 R in the radius distribution of small planets, which separates super-Earths and sub-Neptunes (Owen & Wu, 2013; Fulton et al., 2017). In this work, we used the radius valley for classification because super-Earths and sub-Neptunes may have different formation mechanisms and formation locations (e.g., Chen et al., 2024). To facilitate a more detailed analysis, we calculated the radius valley using
@@
which was derived from Eq. (12) in Zhu & Dong (2021). Based on this, we divided all planetary systems in our sample into three categories: systems containing only super-Earths (Rp $<$ Rgap), systems containing only sub-Neptunes (Rp $\geq$ Rgap), and mixed systems that include both super-Earths and sub-Neptunes within the same system. These three categories represent completely distinct populations with no overlap.
Figure 2 illustrates the architecture of the three population samples of all multis analyzed in this study. The left panel represents super-Earth systems, the middle panel represents sub-Neptune systems, and the right panel represents mixed systems containing both super-Earths and sub-Neptunes within the same system. Each panel is sorted by stellar mass. A trend is evident: As stellar mass increases from low to high, the semimajor axis shifts progressively from smaller to larger values.
In the final sample, we used the following datasets: all multiple systems, consisting of 166 systems with 408 planets; super-Earth systems, including 55 systems with 132 planets; sub-Neptune systems, containing 33 systems with 70 planets; and mixed multiple systems, comprising 78 systems with 206 planets. The “all multiple systems” dataset includes the total data from the other three populations.